OpenStudy (anonymous):

How many integers between 100 and 1000 are multiples of 7? show all work

OpenStudy (anonymous):

hmm.. not sure.... but i think it's ... 1000 ÷ 7 = 142R6 <== means there are 142 multiples of 7 from 1 to 1000 100 ÷ 7 = 14R2 <== means there are 14 multiples of 7 from 1 to 100 so from 100 to 1000, there are 142-14=128 multiples of 7